OpenStudy (anonymous):

The strength of an electric field at 5.0 cm from a point charge is 100.0 N/C. What is the magnitude of the source charge? Please explain your answer.

OpenStudy (irishboy123):

Coulumbs law, force between 2 charges, q1, q2, separated by distance r: F = k q1 q2 /r^2 the field strength E is measured as F/q, force per charge, so the field around a charge q, at distance r from charge, is: E = k q /r^2

OpenStudy (anonymous):

Thank you!

OpenStudy (shamim):

Can u tell me charge q=?

OpenStudy (shamim):

Can yu tell me k=?

OpenStudy (anonymous):

E = k q /r^2 in given problem E= 100.0 N/C r=5.00 cm= 0.05 m k=9x10^9 so equation becomes 100 = 9 x 10^9x q/ 0.05^2 we get q= 2.7x10^-11 C

OpenStudy (anonymous):

k = 4 = 9*10^9 q = 100*5*5/10000*9*10^9 = 2.7 * 10^(-11)